OpenStudy (anonymous):

how do you evaluate ln(e)^1/5 or ln 5root of e

ganeshie8 (ganeshie8):

use the log property : \(\large \ln a^b = b \times \ln a\)

OpenStudy (anonymous):

Oh okay hold on let me see what i get.

ganeshie8 (ganeshie8):

and keep in mind that the base of of \(\ln\) is \(e\)

OpenStudy (anonymous):

\[\log_b(b^x)=x\]

OpenStudy (anonymous):

so would it be lne^1/5=1/5*lne and then solve ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):